# DESY-25-105

### v1

@@ Acknowledgements
We thank Tomoyuki Arakawa, Heeyeon Kim, Lorenzo Mansi, and Tomas Prochazka for useful discussions. We thank the Simons Center for Geometry and Physics and the organizers of the Supersymmetric Quantum Field Theories, Vertex Operator Algebras, and Geometry program, where this work was conceived. M. J. K. is supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award Numbers DE-SC0013528 and QuantISED Award DE-SC0020360. C.L. acknowledges support from DESY (Hamburg, Germany), a member of the Helmholtz Association HGF; C.L. also acknowledges the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy - EXC 2121 “Quantum Universe” - 390833306 and the Collaborative Research Center - SFB 1624 “Higher Structures, Moduli Spaces, and Integrability” - 506632645. The work of J.S. is supported by the National Research Foundation of Korea (NRF) grants RS-2023-00208602 and RS-2024-00405629, and also by the KAIST-KIAS collaboration program.

### v2

@@ Acknowledgements
We thank Tomoyuki Arakawa, Heeyeon Kim, Lorenzo Mansi, and Tomas Prochazka for useful discussions. We also thank Jacques Distler for comments on a previous version of this paper. We thank the Simons Center for Geometry and Physics and the organizers of the Supersymmetric Quantum Field Theories, Vertex Operator Algebras, and Geometry program, where this work was conceived. M. J. K. is supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award Numbers DE-SC0013528 and QuantISED Award DE-SC0020360. M. J. K. is also supported by the Start-up Research Grant for new faculty provided by Texas A&M University. C.L. acknowledges support from DESY (Hamburg, Germany), a member of the Helmholtz Association HGF; C.L. also acknowledges the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy - EXC 2121 “Quantum Universe” - 390833306 and the Collaborative Research Center - SFB 1624 “Higher Structures, Moduli Spaces, and Integrability” - 506632645. The work of J.S. is supported by the National Research Foundation of Korea (NRF) grants RS-2023-00208602 and RS-2024-00405629, and also by the KAIST-KIAS collaboration program.